OpenStudy (anonymous):

simplify (sin4pi/3) - (sin3pi/4) / 4pi/3 - 3pi/4

OpenStudy (anonymous):

Hint: \[Sin \frac{4\prod }{3} = sin (2\prod -\frac{2\prod }{3})\]

OpenStudy (anonymous):

\[sin(2\prod - \Theta ) =- sin\Theta \]